OpenStudy (anonymous):

PLEASE HELP!!! What is the 41st term of the arithmetic sequence where a1 = 17 and a15 = -39 ?

OpenStudy (anonymous):

a) -147 b) -143 c) -139 d) -135

OpenStudy (anonymous):

# of places between the first and last divided by the distance between the first and last values. Then multiply by the # of places to the place value that you are looking for.

OpenStudy (texaschic101):

a15 - a1 = 14d ....-17 - 39 = -56....14d = -56, d = -4 an = a1 + (n - 1)d a41 = 17 + (41 - 1) * -4 a41 = 17 + 40 * -4 a41 = 17 -160 a41 = -143 I think :)

OpenStudy (anonymous):

thanks!

OpenStudy (texaschic101):

YW :)